Land worth $800,000 is purchased by a down payment of $120,000 and the balance in equal monthly installments for 15 years. If interest is at j12 = 4%, determine the buyer’s and seller’s equity in the land at the end of 10 years
Answer #1

Loan Amount = 800,000 - 120,000 = $680,000

Time Period = 15 years

Interest Rate = 4%

Calculating Monthly Payment,

Using TVM Calculation,

PMT = [PV = 680,000, FV = 0, N = 180, I = 0.04/12]

PMT = $5,029.88

Calculating Loan Balance at the end of Year 10,

Using TVM Calculation,

FV = [PV = 680,000, PMT = -5,029.88, N = 120, I = 0.04/12]

FV = $273,117.38

Buyer's Equity = 800,000 - 273,117.38

Buyer's Equity = $526,882.62

Seller's Equity = $273,117.38